OpenStudy (anonymous):

solve by creating same bases: 1/9 = 27 ^ x-1

OpenStudy (mertsj):

\[\frac{1}{9}=3^{-2}\]

OpenStudy (mertsj):

\[27^{x-1}=(3^{3})^{x-1}\]

OpenStudy (mertsj):

\[(3^{3})^{x-1}=3^{3x-3}\]

OpenStudy (mertsj):

So:\[3^{-2}=3^{3x-3}\]

OpenStudy (mertsj):

And -2=3x-3 1=3x 1/3=x

OpenStudy (anonymous):

Can you please help me with one more??? PLEASE!! begging you.

OpenStudy (mertsj):

Post it.

OpenStudy (anonymous):

properties of logarithms: 3^x=5^x+2

OpenStudy (anonymous):

3^x=5^(x+2)

OpenStudy (mertsj):

\[x \log_{10}3=(x+2)\log_{10}5 \]

OpenStudy (anonymous):

How did you get that?

OpenStudy (mertsj):

Find the log of 3 and the log of 5 using your calculator and solve the equation.

OpenStudy (mertsj):

\[a ^{n}=n \log_{10}a \]

OpenStudy (anonymous):

how did you know what the variables were?

OpenStudy (mertsj):

That is the property of law of logs. Perhaps your book uses different variables. What the variables are doesn't matter. Did you book say: \[m ^{n}=n \log_{10}m \]

OpenStudy (anonymous):

oh eyes!

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So use that property on 3^x

OpenStudy (mertsj):

Would you not get xlog3?

OpenStudy (anonymous):

I think my answer is wrong :/

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

4.52

OpenStudy (mertsj):

Hang on.

OpenStudy (mertsj):

xlog3=.477x and (x+2)log 5 = .699x+1.40

OpenStudy (mertsj):

do you agree with that/

OpenStudy (anonymous):

Yes!

OpenStudy (mertsj):

Subtract .699x from both sides.

OpenStudy (mertsj):

-.222x=1.40

OpenStudy (anonymous):

Oh so the final answer is 1.40?

OpenStudy (mertsj):

Now divide both sides by -.222

OpenStudy (anonymous):

Oh so -6.31?

OpenStudy (mertsj):

yes.

OpenStudy (anonymous):

THANK YOU SO MUCH! seriously, you are my life-saver. Now I have one more, I don't want to make you have to show me but could I just ask you a question about it?

OpenStudy (mertsj):

Yes

OpenStudy (anonymous):

http://gyazo.com/8020e5d281ab10e0d5802e2690ad9794

OpenStudy (anonymous):

I don't get if there's acertain equation I am supposed to apply because it looks so confusing.

OpenStudy (mertsj):

This involves two laws of logs. On the left side you have the difference of two logs which is equal to the log of the quotient.

OpenStudy (mertsj):

\[\log_{3} \frac{x+1}{x}\]

OpenStudy (mertsj):

The right side is the property we used in the last problem.

OpenStudy (mertsj):

\[2\log_{3}2=\log_{3}2^{2} \]

OpenStudy (mertsj):

So now, if the logs are equal, and the bases are the same, the arguments must be equal

OpenStudy (mertsj):

\[\frac{x+1}{x}=2^{2}\]

OpenStudy (mertsj):

solve that equation.

OpenStudy (mertsj):

Cross multiply

OpenStudy (mertsj):

4x=x+1

OpenStudy (mertsj):

3x=1 x=1/3

OpenStudy (anonymous):

thank you so so so much!